OpenStudy (anonymous):

Find a polynomial with integer coefficients that satisfies the given conditions. Q has degree 3 and zeros −6 and 1 + i.

OpenStudy (anonymous):

who can solve it?

OpenStudy (anonymous):

Actually, aren't there a family of polynomials that satisfy these conditions? For example if I found a solution, what would stop me from multiplying the whole thing by 2?

OpenStudy (anonymous):

Anyway, we know one factor is (x+6), because ((-6)+6) = 0.

OpenStudy (anonymous):

If 1+i is a root, then so is 1-i. We thus make factors of them: \[ (x - (1+i))(x-(1-i))(x+6) = 0 \]

OpenStudy (anonymous):

It's just a matter of expansion... as I suspected, there are a family of solutions.

OpenStudy (anonymous):

\[ \begin{split} (x - (1+i))(x-(1-i))(x+6) &= (x^2-x-ix-x+1+i+ix-i-i^2)(x+6) \\ &= (x^2-2x+1-i^2)(x+6) \\ &= (x^2-2x+2)(x+6) \\ &= x^3-2x^2+2x + 6x^2 - 12 x + 12 \\ &= x^3+4x^2-10x+12 \end{split} \]So my final answer is \[ \Large x^3+4x^2-10x+12 \]

OpenStudy (anonymous):

But the complete solution set would be \[ \Large \forall c\in \mathbb{Z} \quad cx^3+4cx^2-10cx+12c \]Since any integer coefficients will fly.

OpenStudy (anonymous):

thank you WO

OpenStudy (anonymous):

WIO*